OpenStudy (anonymous):

The distance, d, an object travels is represented by the formula d = rt, where r is the rate at which the object is traveling and t is the duration of time over which the object travels. If d has units of meters and r has units of meters per minute, what must be the units of t? A. minutes B. meters per second C. seconds D. meters

OpenStudy (triciaal):

d = r * t t = d/r = meters divided by meters per min when you divide by a fraction invert and multiply